Montserrat Volcano Observatory, Montserrat, West Indies

Scientific Report 56
08 March 1997

This report covers the 7-day period to 24:00 on 8 March.


Activity at the volcano has continued to be dominated by intermittent earthquake swarms, rockfall activity and minor pyroclastic flows. Growth of the dome complex within English's Crater is continuing.

Visual Observations

Good views of the volcano have been obtained throughout the week. All observations indicate that the dome complex is continuing to grow. Slow upward growth of the January 20 dome continues but at a slower rate than observed previously. The highest point on the dome was measured at 942 metres on March 3rd.

The eastern face of the dome complex is still steep and continues to be unstable. The north-eastern face of the October dome is beginning to show signs of activity. New rockfall chutes have developed on this face. At the end of the week the whole of this area of the dome was crumbling.

Theodolite measurements of points on the north and north-eastern faces of the October dome show that there is some swelling occurring in this area. Rates of outward movement of this area have been of the order of 1.5 m/day from 1-7th March.

Night-time observation of the dome complex revealed strong incandescence from the active areas of the January 20 dome and in the chutes on the October dome. The incandescent area changed throughout the week from the eastern face to the northern and north-eastern faces. Faint incandescence was observed over most of the October dome.

Small rockfalls have occurred almost continuously throughout the reporting period sourcing from the active faces of the January and October domes. Some of these developed into small pyroclastic flows but none of these travelled any further than the track crossing the Tar River Valley. Small ash clouds associated with these events have drifted westward on the prevailing winds.

Galway's Wall has shown little outward signs of change. Slow disintegration is continuing. Only one small rockfall was observed which travelled part-way down the crater wall.


The level of seismicity has been relatively low for the major part of the reporting period. Two earthquake swarms early in the week were dominated by hybrid events. The remainder of the week has been quiet with the main source of seismicity being the rockfalls and small pyroclastic flows.

Table 1: Earthquake types
These earthquake counts are of events that triggered the short-period seismic network event recording system between 0000 and 2400 each day.

Date	VT	LP	Hybrid	Dome RF
02 MAR 97	21	0	91	21
03 MAR 97	2	0	2	38
04 MAR 97	23	1	47	24
05 MAR 97	2	0	0	33
06 MAR 97	0	0	1	11
07 MAR 97	0	1	1	9
08 MAR 97	1	0	2	8

Table 2: VT / hybrid earthquake swarms and tremor episodes

Date	Start time	Durn (hrs)	VTs	Hybrid
01 MAR 97	21:54	17.95	21	95
04 MAR 97	03:27	12.75	20	43

Ground Deformation

EDM measurements were made on radial lines from the volcano in the western northern triangles and southern. The line from Upper to Lower Amersham showed a shortening of 21mm while the line from Lower Amersham to Chances Peak Steps lengthened by 22 mm over the period since February 19. This continues the long-term trend of overall stability of these lines.

Lines from Windy Hill to Farrells and St George's Hill to Farrells changed by +4mm and -2mm, respectively. These measurements continue the long-term trends of stability of these lines.

The line from Galway's Plantation to Chances Peak was re-measured during the week following the cleaning of ash from the reflector. This line had shortened by 15mm since it was last measured on February 17. This continues the slow shortening trend seen on this line. The total shortening is now of the order of 10cm over the past 13 months.

Occupations of the GPS networks, WESTNET and EASTNET were carried out on the 5th and 6th of March respectively. The absence of any significant changes in baseline length, station position or station height indicate no the absence of a widespread strain field associated with the eruption.

GPS occupations of a site (FT3) 80m NW of the dome have shown large movements. The baseline to Harris (M18) shortened by 6.5cm between January 18th and March 3rd. Movement of the station is radially away from the dome (i.e. to the NW) by 12.9cm (March 3rd) and 14.5cm (March 8th) since January 18th. The station has shown no significant change in height, all occupations giving a height within 1.6cm of one another.

Crack 2 on Chances Peak was measured on March 1st and 8th. The opening rate of the crack is constant at around 0.5 mm per day. Differential movement between the nails indicates that the crack has broken up into a zone of fracturing rather than a discrete crack. Shear displacement along the crack which began at the end of January has now reached 18cm total movement.

Dome Volume Measurements

No dome volume surveys were carried out during the reporting period. However the last well-established dome volume on March 1st is 42 x 106 m3. The total erupted magma is estimated at 72 x 106 m3, including pyroclastic flow deposits (expressed as dense rock equivalent).

Gas Measurements

Measurements of sulphur dioxide flux were made using the MiniCOSPEC on 2 March (106 t/d), 3 March (174 t/d), 5 March (159 t/d), 6 March and 7 March. The fluxes from the latter two dates are yet to be calculated but the data show low tenor plume traces consistent with low values.

Environmental Monitoring

Rain water samples were collected at 4 locations around the volcano on 3 March, and results are shown in Table 3. One sample was collected from the overflow to a water storage tank at Fairfield (Trials Reservoir).

These results show that the rainwater directly west of the volcano continues to be highly acidic and has high concentrations of certain anions, although the pH has increased slightly. The water from the storage tank has anion concentrations and pH levels well within WHO guidelines for drinking water.

Table 3: Rain and surface water geochemistry, 2 Mar 97
Units: conductivity (mS/cm), total dissolved solids (g/l), sulphate (mg/l), chloride (mg/l), fluoride (mg/l),

			pH      Cond.   TDS     Slphts  Chlrds  Flrds
Upper Amersham		2.60	0.448	0.224	24	118	>1.5
Lower Amersham		2.85	0.911	0.456	3	52	1.4
Police HQ, Plymouth	3.31	0.237	0.118	3	53	0.8
Weekes			6.20	0.402	0.201	3	44	0.65
Trials Reservoir	7.68	0.618	0.309	38	90	0.35

Ash was collected at two sites to the west of the volcano, but appeared to be mostly wind-blown. No further results on ash analyses were received during this period.

Staff Changes
Mr Godfrey Almorales (SRU)
Dr Paul Jackson (SRU)
Prof Steve Sparks (Bristol University, UK)
Dr Paul Cole (Luton University, UK)

Prof Ramsay Saunders (SRU), Dr Tony Reedman (BGS) & Mr Lloyd Lynch (SRU) were briefly on island to attend the first MVO Management Board meeting held on Friday, 7th March.

Montserrat Volcano Observatory